Green plants use light from the Sun to drive photosynthesis. Photosynthesis is a chemical reaction in which water H 2 O and carbon dioxide CO 2 chemically react to form the simple sugar glucose C 6 H 12 O 6 and oxygen gas O 2 . What mass of water is consumed by the reaction of 3.04 g of carbon dioxide? Round your answer to 3 significant digits.
Answer #1

Ans :

The reaction is given as :

6CO2 + 6H2O = C6H12O6 + 6O2

Number of mol of CO2 = 3.04 / 44.01 = 0.069 mol

6 mol CO2 consumes 6 mol of water

So 0.069 mol of CO2 will consume 0.069 mol of water

Then the mass of water consumed = 0.069 x molar mass

= 0.069 x 18.01528

= 1.24 grams